The base of a triangle exceeds the height by 5m, If the area is 42 square meters, find the length of the base and the height of the triangle.
Answer by Alan3354(69443)

The base of a triangle exceeds the height by 5m, If the area is 42 square meters, find the length of the base and the height of the triangle.
-----------
Area = b*h/2 = 42
b*h = 84
-------
Find a pair of factors of 84 that differ by 5.